OpenStudy (ccastudentgrade10):

HELP PLS FAN AND MEDAL A gas has a volume of 675 mL at a temperature of –45.0°C. What volume will the gas occupy at 25.0°C? Show your work.( points awarded 1 point for correct answer, 1 point for work, 1 point for correct set up ) I missed chem class and i have to finish this by 11, there are seven more of these problems too but i just need thsi one explained and i can do the rest

OpenStudy (aaronq):

Assuming ideal gas behaviour, we can use the ideal gas law - a combination of charles' law, boyle's law, etc. In a closed system, the physical parameters of P, V and T have to remain constant, meaning that if we have a system at equilibrium and we change one (or more) of the parameters, the other parameters will have to change as well to make this equality true: \(\sf system_1=system_2\) \(\sf \dfrac{P_1V_1}{nRT_1}=\dfrac{P_2V_2}{nRT_2}\) since \(n\) and \(R\) are constants we can omit them, leaving us with: \(\sf \dfrac{P_1V_1}{T_1}=\dfrac{P_2V_2}{T_2}\) This formula can solve all these problems, for example, solving for volume here: \(\sf V_2=\dfrac{(675~mL)}{(-45+273)K}*(25+273)K\) (we ignore P because it doesn't change, and they it will just cancel itself to 1).

OpenStudy (ccastudentgrade10):

is the answer 882.08 then?

OpenStudy (ccastudentgrade10):

@aaronq

OpenStudy (aaronq):

you need the right units, but the numerical value is right

OpenStudy (ccastudentgrade10):

so 882.08K ?

OpenStudy (aaronq):

no, K is a unit of temperature. If you carry out the dimensional analysis through, you'll the units are "mL"

OpenStudy (ccastudentgrade10):

so 882.08ml then?

OpenStudy (aaronq):

yep

OpenStudy (ccastudentgrade10):

okay thank you!

OpenStudy (aaronq):

no problem